# The Resolution of the Gauge Problem of Cosmology provides Insight into the Formation of the First Structures in the Universe

### v2

@@ Appendix C Conventional Evolution Equation for Cosmological Density Fluctuations
where Eq. (71b) has been used.
In order to arrive at a generalization of (150), the following steps must be taken. First, it is necessary to differentiate equation (151a) with respect to time. Subsequently, the time derivatives of $w$, $H$, $\varepsilon_{\scriptscriptstyle(0)}$, $\vartheta_{\scriptscriptstyle(1)}$ and ${{}^{3}\!}R_{\scriptscriptstyle(1)}$ must be eliminated. This can be achieved by utilizing the equations (44), (71a) and (71b) and the perturbation equations (151b) and (151c). The final step in the process is the elimination of ${{}^{3}\!}R_{\scriptscriptstyle(1)}$ through the application of (151a). This results in the following system of equations:
@@
$\qquad-\biggl{[}\beta^{2}\dfrac{\nabla^{2}}{a^{2}}+\tfrac{1}{2}\kappa\varepsilon_{\scriptscriptstyle(0)}\Bigl{(}(1+w)(1+3w)\Bigr{.}\Bigr{.}$
@@
In this procedure the computer algebra program Maxima [44] has been used.
The gauge modes (147) are for the system (152) given by
@@
The fact that the gauge modes (153) satisfy the system (152) has been shown in Appendix B.
It can be inferred from the set of equations (152) and the gauge modes (153) that:
  • The solution of the homogeneous part of Eq. (152a) is a linear combination of a physical mode and the gauge mode (153a). The particular solution is obtained by solving Eq. (152b), and contains the gauge mode (153b).
  • It is impossible to determine the arbitrary infinitesimal gauge function $\psi(\bm{x})$ through physical means. Therefore, the imposition of physical initial values on the system (152) does not yield a physical solution.
  • In the small-scale limit, the gauge modes (153) do not vanish. This means that in the small-scale limit gauge-dependent quantities do not become gauge-invariant. Consequently, $\delta:=\varepsilon_{{\scriptscriptstyle(1)}}/\varepsilon_{{\scriptscriptstyle(0)}}$ does not become equal to its Newtonian counterpart $\delta_{\varepsilon}:=\varepsilon^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}/\varepsilon_{{\scriptscriptstyle(0)}}$, given by (56) and (58). This is consistent with the gauge transformation (53) of the Newtonian theory of gravitation.
  • A notable distinction between the conventional equation (150) and Eq. (152a) is that the former is homogeneous, whereas Eq. (152a) is inhomogeneous. This is because the system (152) is derived from the complete set of constraint equations and conservation laws. In the case where the right-hand side of Eq. (152a) is absent, as is the case in Eq. (150), then (153a) is a gauge mode if $\nabla^{2}\psi=0$. Consequently, $\psi=C$ is an arbitrary infinitesimal constant, so that the gauge mode (153a) does not vanish.
  • In the universe following decoupling, the values of $w\ll 1$ and $\beta^{2}\ll 1$ are negligible, allowing for the subsequent neglect of $w$ in relation to the constants of order one in Eq. (150). This is equivalent to deriving the usual equation (150) from the Newtonian theory of gravitation adapted to the expansion of the universe. See, for example, Ref. [10], Sec. 6.2, Ref. [31], Sec. 15.9, and Ref. [45], Eq. (3.17). However, the gauge transformation (53) in the nonrelativistic limit ensures that the gauge mode (153a) does not vanish.
From these points it is clear that the conventional equation (150) is unsuitable to study the evolution of density perturbations in the universe. Given that Eq. (150) can be derived from Newtonian gravitation for $w\ll 1$, it follows that the Newtonian theory is unable to provide an adequate description of density perturbations in the universe.
In the ensuing two sections, a comparison is presented between Eqs. (79) and (95), with the system of equations (152). This comparison will be conducted in the case of large-scale fluctuations. It will become evident that the gauge modes preclude the attainment of physical solutions.
@@ C.1 Radiation-Dominated Era
In this era, the pressure is given by a linear barotropic equation of state $p=\tfrac{1}{3}\varepsilon$, so that $\beta^{2}=\tfrac{1}{3}$ and ${w=\tfrac{1}{3}}$. In this case Eqs. (152) result in the system
@@
The gauge modes (153) are solutions of the system (154) for $w=\tfrac{1}{3}$, as can be verified by substitution.
For large-scale fluctuations $\nabla^{2}\delta^{\textnormal{phys}}\rightarrow 0$, the solution of the system (154) is
@@
where it is used that $H=1/2t$ and $\kappa\varepsilon_{\scriptscriptstyle(0)}=3/4t^{2}$. The expression between parentheses in (155a) is the solution of the homogeneous part of Eq. (154a). The particular solution $\tfrac{9}{8}t^{1/2}$ is a consequence of $\vartheta^{\textnormal{phys}}_{\scriptscriptstyle(1)}=\tfrac{9}{8}t^{-1/2}$. The solution (85) derived from the novel perturbation equation (82) is not derivable from the standard equations (154) due to the presence of the gauge function $\psi$ in the solution (155a).
Since $\nabla^{2}\delta^{\textnormal{phys}}$ could have been large for small-scale fluctuations, it may have had a large influence on $\vartheta^{{\textnormal{phys}}}_{\scriptscriptstyle(1)}$ and this could have, in turn, a major impact on the evolution of $\delta^{\textnormal{phys}}$. That is why (79a) yields oscillating density fluctuations (84) with an increasing amplitude, instead of a constant amplitude which follows from (150).
@@ C.2 Era after Decoupling of Matter and Radiation
In this era, the equation of state for the pressure is according to thermodynamics given by (88). Therefore, (91) and (92) imply that $w\approx\tfrac{3}{5}\beta^{2}\ll 1$, so that with $T_{\scriptscriptstyle(0)}\propto a^{-2}$ one obtains $\dot{\beta}/\beta=-H$. Using Eq. (71a) one can derive the following result: $6\beta\dot{\beta}H=-2\kappa\varepsilon_{\scriptscriptstyle(0)}\beta^{2}$. Upon substituting the latter expression into (152a) and neglecting $w$ and $\beta^{2}$ with respect to the constants of order one, the system (152) results in
@@
The gauge modes (153) are solutions of the system (156) for $w\ll 1$ and $\nabla^{2}\psi=0$, as can be verified by substitution. Consequently, for the system (156) $\psi$ is an arbitrary infinitesimal constant $C$. This implies that $\vartheta_{\scriptscriptstyle(1)}=\vartheta^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ is a physical quantity, since its gauge mode $\hat{\vartheta}_{\scriptscriptstyle(1)}$, (153), vanishes identically. Consequently, for large-scale fluctuations the solution of the system (156) is
@@
where $H=2/3t$, $\kappa\varepsilon_{\scriptscriptstyle(0)}=4/3t^{2}$, and $\beta^{2}\propto t^{-4/3}$ has been used. The latter proportionality follows from (91) and (92), $T_{\scriptscriptstyle(0)}\propto a^{-2}$ and (104). The expression between parentheses in (157a) is the solution of the homogeneous part of Eq. (156a). The particular solution $\tfrac{7}{9}t^{-5/3}$ is a consequence of $\vartheta^{\textnormal{phys}}_{\scriptscriptstyle(1)}=-\tfrac{7}{9}t^{-4/3}$. The solution (109) derived from the novel perturbation equation (108) is not derivable from the standard equations (156) due to the presence of the gauge constant $C$ in the solution (157a).
Since $\psi=C$, the gauge transformation (1) with $\xi^{\mu}(t,\bm{x})$ given by (6) reduces to the Newtonian gauge transformation (53). This is to be expected, since a cosmological fluid for which ${w\ll 1}$ and $\beta^{2}\ll 1$ can be described by nonrelativistic equations of state (88).

### v3

@@ Appendix C Conventional Evolution Equation for Relative Density Perturbations
where Eq. (75b) has been used. First, differentiate equation (150a) with respect to time. Then, eliminate the time derivatives of $w$, $H$, $\varepsilon_{\scriptscriptstyle(0)}$, $\vartheta_{\scriptscriptstyle(1)}$ and ${{}^{3}\!}R_{\scriptscriptstyle(1)}$ using equations (44), (75a) and (75b) and the perturbation equations (150b) and (150c). Finally, eliminate ${{}^{3}\!}R_{\scriptscriptstyle(1)}$ using Eq. (150a). This results in the system of equations:
@@
[…]
@@
In this procedure the computer algebra program Maxima [52] has been used. The gauge modes (147) are for the system (151) given by
@@
From the Friedmann equation, $3H^{2}=\kappa\varepsilon_{\scriptscriptstyle(0)}$, it follows that $H\neq 0$ for $\varepsilon_{{\scriptscriptstyle(0)}}\neq 0$. This implies that the gauge mode $\hat{\delta}(t,\bm{x})$ is nonzero, regardless of the scale of the density perturbation. Consequently, the general solution of the system (151) has no physical meaning. Therefore, the system (151) cannot be used to study the evolution of density perturbations.
@@ C.1 Radiation-Dominated Era
In this era, the pressure is given by a linear barotropic relativistic equation of state $p=\tfrac{1}{3}\varepsilon$, so that $\beta^{2}=\tfrac{1}{3}$ and $w=\tfrac{1}{3}$}. The standard equation is given by
$\ddot{\delta}+2H\dot{\delta}-\left(\dfrac{1}{3}\frac{\nabla^{2}}{a^{2}}+\tfrac{4}{3}\kappa\varepsilon_{\scriptscriptstyle(0)}\right)\delta=0.$ (153)
In contrast, Eqs. (151) result in the system
@@
The gauge modes (152) are solutions to the system (154) for $w=\tfrac{1}{3}$, as can be verified by substitution.
For large-scale perturbations $\nabla^{2}\delta^{\textnormal{phys}}\rightarrow 0$, the solution to the system (154) is
@@
where it is used that $H=\tfrac{1}{2}t^{-1}$ and $\kappa\varepsilon_{\scriptscriptstyle(0)}=\tfrac{3}{4}t^{-2}$. The expression between parentheses in (155a) is the solution to the homogeneous part of Eq. (154a). The particular solution $\tfrac{9}{8}t^{1/2}$ is a consequence of $\vartheta^{\textnormal{phys}}_{\scriptscriptstyle(1)}=\tfrac{9}{8}t^{-1/2}$. The right-hand side of Eq. (154a) is not equal to zero, which means that Eq. (153) is incomplete. As a result, it does not provide all physical solutions. Since $\nabla^{2}\delta^{\textnormal{phys}}$ could have been large for small-scale perturbations, it may have had a large influence on $\vartheta^{{\textnormal{phys}}}_{\scriptscriptstyle(1)}$ and this could have, in turn, a major impact on the evolution of $\delta^{\textnormal{phys}}$. Because the quantity $\vartheta_{{\scriptscriptstyle(1)}}$ is incorporated into Eq. (82), oscillating relative density perturbations (87) with an increasing amplitude are obtained instead of the constant amplitude that follows from (153).
@@ C.2 Era after Decoupling of Matter and Radiation
The standard equation is given by
$\ddot{\delta}+2H\dot{\delta}-\left(\beta^{2}\dfrac{\nabla^{2}}{a^{2}}+\tfrac{1}{2}\kappa\varepsilon_{\scriptscriptstyle(0)}\right)\delta=0,$ (156)
where $\beta$ is given by (93). In this era, the equation of state for the pressure is according to thermodynamics given by (89). Therefore, (92) and (93) imply that $w\approx\tfrac{3}{5}\beta^{2}\ll 1$, so that with $T_{\scriptscriptstyle(0)}\propto a^{-2}$ one obtains $\dot{\beta}/\beta=-H$. Using Eq. (75a) one can derive the following result: $6\beta\dot{\beta}H=-2\kappa\varepsilon_{\scriptscriptstyle(0)}\beta^{2}$. Upon substituting the latter expression into (151a) and neglecting $w$ and $\beta^{2}$ with respect to the constants of order one, the system (151) results in
@@
The gauge modes (152) are solutions to the system (157) for $w\ll 1$ and $\nabla^{2}\psi=0$, as can be verified by substitution. Consequently, for the system (157) $\psi$ is an arbitrary infinitesimal constant $C$. This implies that $\vartheta_{\scriptscriptstyle(1)}=\vartheta^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ is a physical quantity, since its gauge mode $\hat{\vartheta}_{\scriptscriptstyle(1)}$, (152b), vanishes identically. Consequently, the relativistic gauge transformation (3) with $\xi^{\mu}(t,\bm{x})$ given by (9) reduces to the Newtonian gauge transformation (52). This is to be expected, since a cosmological fluid for which $w\ll 1$} and $\beta^{2}\ll 1$ can be described by nonrelativistic equations of state (89).
For large-scale perturbations $\nabla^{2}\delta^{\textnormal{phys}}\rightarrow 0$, the solution to the system (157) is
@@
where $H=\tfrac{2}{3}t^{-1}$, $\kappa\varepsilon_{\scriptscriptstyle(0)}=\tfrac{4}{3}t^{-2}$, and $\beta^{2}\propto t^{-4/3}$ has been used. The latter proportionality follows from (92) and (93), $T_{\scriptscriptstyle(0)}\propto a^{-2}$, (94), and $a\propto t^{2/3}$, (104). The expression between parentheses in (158a) is the solution to the homogeneous part of Eq. (157a). The particular solution $\tfrac{7}{9}t^{-5/3}$ is a consequence of $\vartheta^{\textnormal{phys}}_{\scriptscriptstyle(1)}=-\tfrac{7}{9}t^{-4/3}$.
The right-hand side of Eq. (157a) is nonzero for all scales. Therefore, Eq. (156) is incomplete and does not yield all physical solutions.